OpenStudy (anonymous):

The number of real roots of the equation , (x-1)^2 + (x-2)^2 + (x-3)^2=0

OpenStudy (anonymous):

I exapanded the backets and got the answer , but i am sure there is a better way , so i have asked this poblem here

OpenStudy (ikram002p):

its quadratic equation at the end right ? hmm so u got 2 roots ?

OpenStudy (cwrw238):

yep

ganeshie8 (ganeshie8):

yeah equation is missing

OpenStudy (anonymous):

No real roots

OpenStudy (anonymous):

Which equation is missing?

ganeshie8 (ganeshie8):

I don't see an equation above, are we to assume that expression = 0 ?

OpenStudy (cwrw238):

yes thats what i assumed

OpenStudy (cwrw238):

should have asked really...

ganeshie8 (ganeshie8):

a square can never be negative, so if the equation has a real solution, then it should evaluate the value of expression to 0. hmm still thinking...

OpenStudy (ikram002p):

(x-1)^2 + (x-2)^2 + (x-3)^2 >=0 lets assume it zero then the only way x=1 &x=2&x=3 (which is a contradiction ) thus no real else expand and see hmmm

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

I didn't get ikram , sorry it is equal to zero , i forgot to put @ikram002p

OpenStudy (anonymous):

@ganeshie8 @ikram002p

OpenStudy (ikram002p):

what i mean ur equation is :- (x-1)^2 + (x-2)^2 + (x-3)^2=0 now left side is always positive ok ? (sum of squars ) and each term is also positive thus the only way to get zero is each term =0 xD hope this is help ='(

OpenStudy (anonymous):

SO, (x-1)^2 + (x-49)^2 + (x+5665)^2 +(x-66)^2 =0 can never have real roots??

OpenStudy (anonymous):

with the same logic

OpenStudy (ikram002p):

humm yes

OpenStudy (anonymous):

nice , if you are sure ^_^

OpenStudy (ikram002p):

positive sure :P

OpenStudy (anonymous):

the logic is good , thank you